OpenStudy (anonymous):

integral of sin(3x) cos^2(3x) dx

OpenStudy (anonymous):

The answer is: \[-\frac{1}{9} \cos ^3(3 x) \] Make the change of variable u=cos(3x) du = - 3 sin(3 x) dx and you can make the final step.

OpenStudy (ash2326):

@ncorrea92 did you understand how to solve the problem?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

substitute\[u=\cos \left( 3x \right)\]\[du=-3\sin \left( 3x \right)dx\]\[-\frac{1}{3}du=\sin \left( 3x \right)dx\]

OpenStudy (anonymous):

Thanks i got it

OpenStudy (ash2326):

Okay:) we have \[\int \sin (3x) \cos^2 (3x) dx\] We notice that there is sin 3x and cos 3x , so we can use substitution here, either sin 3x or cos 3x if we substitute sin 3x it won't help much Let's substitute \[\cos 3x = u\] differentiate this with respect to x \[-3 \sin 3x=\frac{du}{dx}\] so \[-3\ \sin 3x\ dx= du\] or \[\sin 3x\ dx =-\frac{du}{3}\] Let's substitute this in the integral \[\int \sin 3x\ \cos^2 3x\ dx\] let's substitute now \[\int -u^2\ \frac{du}{3}\] so we get now \[-\frac{u^3}{3 \times 3} +c\] \[-\frac{u^3}{9}+c\] u= cos 3x so we get \[-\frac{\cos^3 3x}{9}+c\] c= is a constant